OpenStudy (anonymous):

prove cos x+ 1/cot x = sin x + tan x

OpenStudy (anonymous):

confirm \(\dfrac{cos x +1}{cot x}\) or cosx + \(\dfrac{1}{cot x}\)??

OpenStudy (anonymous):

cos x+1/ cot x the first one you have up there

OpenStudy (anonymous):

do you know cot = 1/tan or tan =1/cot?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, just replace 1/cot = tan, the left hand side of your expression is tanx (cosx +1) = tanx cosx + tanx = the right hand side, that's it

OpenStudy (anonymous):

what about the sin

OpenStudy (anonymous):

hey tanx = sin*cos, so tan *cos =sin

OpenStudy (anonymous):

okay thanks